A rectangle whose length is 3cm more than its width has an area of 40 square centimeters. Find the length and the width of this rectangle.
Answer by checkley77(12844)

L=W+3
LW=40
(W+3)W=40
W^2+3W-40=0
(W+8)(W-5)=0
W-5=0
W=5 CM. FOR THE WIDTH
L=5+3=8 CM FOR THE LENGTH.
PROOF:
5*8=40
40=40